OpenStudy (anonymous):

a cylinder has a surface area of 339cm. The radius is two times greater than the height. what is the radius of the cylinder. can you explain the steps to getting the answer aswell thanks.

OpenStudy (campbell_st):

Rewrite the height in terms of r. let the radius be r then the height is r/2 then for a closed cylinder \[SA = 2 \pi r^2 + 2 \pi r h\] \[339 = 2\pi r^2 + 2\pi r (r/2)\] gives \[339 = 2\pi r^2 + \pi r^2\] \[339 = 3\pi r^2\] \[113/\pi = r^2\] \[r = \sqrt{113/\pi}\]

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

how to substact a fraction?

hero (hero):

substract or subtract?